OpenStudy (anonymous):

Find the inclination, theta, of the line passing through the points (5,0) and (-5,5)

OpenStudy (anonymous):

could someone tell me the equation to use?

OpenStudy (rajee_sam):

Inclination is actually the slope which is \[\tan \theta \] So find the slope first using the points given. Slope = \[\frac{ y _{2} - y _{1}}{ x _{2} - x _{1} }\]

OpenStudy (rajee_sam):

Can you find the slope?

OpenStudy (anonymous):

it is -.5

OpenStudy (rajee_sam):

yes

OpenStudy (rajee_sam):

now \[\tan \theta = -0.5\] So \[\theta = \tan^{-1} (-0.5)\] use calculator to find it

OpenStudy (anonymous):

it gives me - 26.6

OpenStudy (anonymous):

i have the choices of 26.6, 153.4, 63.4, 116.6 or none of these

OpenStudy (anonymous):

and so would it 26.6 ?

OpenStudy (anonymous):

oh. okay! Thank you!

OpenStudy (rajee_sam):

My earlier answer is right bad teacher. it is 153.4 and that is final

OpenStudy (rajee_sam):